OpenStudy (anonymous):

Convert the polar equations r=6 Cos 4 θ

ganeshie8 (ganeshie8):

multiply both sides wid \(r\)

ganeshie8 (ganeshie8):

\(r^2 = 6r\cos 4\theta \)

ganeshie8 (ganeshie8):

\(x = r\cos 4\theta\) \(y = r\sin 4\theta\) \(r^2 = x^2 + y^2\)

ganeshie8 (ganeshie8):

the equation becomes : \(x^2 + y^2 = 6x\)

OpenStudy (anonymous):

Thank u

ganeshie8 (ganeshie8):

np :)

ganeshie8 (ganeshie8):

ive got second thoughts on this

ganeshie8 (ganeshie8):

looks i messed it up, @ikram002p help plz

ganeshie8 (ganeshie8):

@Diana97 i think we cannot simpify like that.

ganeshie8 (ganeshie8):

the answer wud be : \(x^2 + y^2 = 6\cos (4*\tan^{-1}(y/x))\)

ganeshie8 (ganeshie8):

it wont simplify i believe.